OpenStudy (anonymous):

solve (4^2)^3 show work

OpenStudy (anonymous):

\(4^2\) = ??

OpenStudy (mayankdevnani):

\[\large \bf {(4^2)}^3\] first of all,what is the value of \[\large \bf 4^2\]

OpenStudy (anonymous):

After finding that "??": \(??^3\) = ?????

OpenStudy (mayankdevnani):

and remember : \[\large \bf x^2=x \times x\]

OpenStudy (anonymous):

That "?????" is your answer.. :P

OpenStudy (mayankdevnani):

@spencerbeymer

OpenStudy (anonymous):

@mayankdevnani you forget the question, he has written : "Show work", so you have to show your work, he will not show anything.. :P

OpenStudy (anonymous):

@mayankdevnani yes that is my question how did you type it out like that?

OpenStudy (anonymous):

With keyboard, I guess..!! :P

OpenStudy (anonymous):

@mayankdevnani 4^2 = 16

OpenStudy (anonymous):

Now, \(16^3\) = ??

OpenStudy (mayankdevnani):

just copy it :- `\[4^2\]`

OpenStudy (mayankdevnani):

@spencerbeymer

OpenStudy (mayankdevnani):

good....